# Tool-Aided Evolutionary LLM for Generative Policy Toward Efficient Resource Management in Wireless Federated Learning

### v1

@@ III Decoupling of the Optimization Problem
Given the feasible device selection action $\mathbb{S}_{t}$, the performance of global model $\varXi$ is independent of the resource allocation action $\mathbb{R}_{t}$, i.e.,
@@
$\mathrm{s.t.}$ $\mathbb{S}_{t}\subset\mathcal{N},$ (20)
@@
The problem decomposition contributes significantly to reducing computational complexity by searching over a smaller-dimensional solution space. On this premise, we can introduce LLM to solve $\mathfrak{P}3$ conveniently, while resorting to a tool-aided solution of $\mathfrak{P}2$. Furthermore, a concerted, evolutionary effort is put forward to adapt to dynamic decision-making environments.
@@ IV-A Tool-Aided LLM for Efficient Policy Generation
where $\mathbf{s}^{\text{text}}_{t}$ represents the added natural language description. This linguistic representation $\mathbf{o}_{t}$, which will be elaborated later, serves as the input to the LLM. Consequently, it efficiently adapts to diverse state spaces across various scenarios without requiring additional tuning of NN architectures. Then, the LLM generates a text response $\tilde{\mathbf{a}}_{t}$ based on the current language description according to its strategy $\pi_{\theta}$, i.e.,
@@
which is interpreted as the decision action for device selection $\mathbb{S}_{t}$. Upon selecting the subset $\mathbb{S}_{t}$ of devices, the LLM invokes an external resource tool to make the resource allocation action $\mathbb{R}_{t}$ for the selected devices. Finally, the joint decisions $\mathbf{a}_{t}=[\mathbb{S}_{t},\mathbb{R}_{t}]$ are applied to the wireless FL environment. Thus, the next observation can be obtained by the environment, i.e.,
@@ IV-B Model-Based Virtual Environment
Due to the significant communication overhead associated with online interactions in wireless FL environments, this section proposes a virtual environment based on an offline model-based approach and tool-assisted computations. Benefiting from the following lemma, the learning of the virtual environment can avoid the extensive use of huge end-to-end expert data.
@@
The state transition of the wireless FL environment can be converted to two parts as:
@@
By definition, $\mathbf{s}^{c}$ and $\mathbf{s}^{m}$ are mutually independent. Therefore, we can have
@@
$P(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})=P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},(\mathbb{S}_{t},\mathbb{R}_{t}))P(\mathbf{s}^{c}_{t+1}|\mathbf{s}^{c}_{t},(\mathbb{S}_{t},\mathbb{R}_{t}))$ (29)
Recalling that by definition, $\mathbf{s}^{m}$ is $\{|D_{n}|,F_{n}(\omega_{n}),\varXi(\mathbb{S}_{t-1})\}$-related. Then, according to the Lemma 1, we have the conclusion. ∎
@@
However, for characterizing the statistics part $P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})$, it remains challenging when dynamically selected devices participate in training and aggregation. To address this, we employ a model to approximate the convergence behavior and performance of the global model during FL training. In other words, simulating the statistics part $P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})$ means to predict $\mathbf{s}^{m}_{t+1}$ including $\varXi(\mathbb{S}_{t})$ based on the current observation $\mathbf{s}^{m}_{t}$, device selection action $\mathbb{S}_{t}$ and the historical accuracy $\varXi(\mathbb{S}_{t-1})$. Generally, $\mathbf{s}^{m}_{t}$ can be classified as the state information pertinent to the selected devices $\mathbf{s}^{m}_{t}(s)$ and that of all devices $\mathbf{s}^{m}_{t}(l)$. In particular, $\mathbf{s}^{m}_{t}(s)$ includes post-training local loss $F_{n}(\bm{\omega}_{n})$, the inner-product $\langle\nabla F_{n}(\bm{\omega}_{n}^{t}),\nabla F_{G}(\bm{\omega}_{G}^{t})\rangle$, and the percentage of the same sign $e(\bm{\omega}_{n}^{t},\bm{\omega}_{G}^{t})$ shared by between local and global models (i.e., $\bm{\omega}_{n}^{t}$ and $\bm{\omega}_{G}^{t})$) for all selected devices $n\in\mathbb{S}_{t}$. Meanwhile, $\mathbf{s}^{m}_{t}(l)$ consists of the local dataset size $|D_{n}|$ (which remains constant during training) and locally computed loss $l_{n}$ of other non-selected devices. We denote the tuple $((\mathbf{s}^{m}_{t},\varXi(\mathbb{S}_{t-1}),\mathbb{S}_{t}),(\mathbf{s}^{m}_{t+1},\varXi(\mathbb{S}_{t})))$ as one sample of the environment model, while several continuous samples constitute a trajectory. The NN architecture of the simulated model is shown in the right part of Fig. 4. During training, we optimize the model with a Mean Squared Error (MSE) loss.
The reward $r_{t}$ can be computed as in (12). Then, alongside the reward $r_{t}$, the virtual environment can respond to $(\mathbf{s}^{m}_{t},\varXi(\mathbb{S}_{t-1}),\mathbb{S}_{t})$, and accurately yield the next state $(\mathbf{s}^{c}_{t+1},\mathbf{s}^{m}_{t+1})$.
@@ IV-C GRPO-Based LLM Training
The aforementioned virtual environment lays the very foundation for updating the LLM agent in an RL manner. Specifically, we employ GRPO [46] to enhance the inference capability of LLM. As a variant of Proximal Policy Optimization (PPO) [49], GRPO also adopts twin policy models, where an old policy $\pi_{\theta_{\text{old}}}$ is used for sampling actions $\tilde{\mathbf{a}}_{t}$ and accumulating transitional records $(\mathbf{o}_{t},\tilde{\mathbf{a}}_{t},r_{t},\mathbf{o}_{t+1})$ in a replay buffer. Here, $r_{t}$ is the reward in the $t$-th round calculated by (12). All samples are described in language, a specific example as shown in Fig. 3. For a batch of $K$-length episodes’ records, where each episode $i$ is denoted as $\{r_{1},\cdots,r_{t},\cdots,r_{K}\}_{i}$, we normalize these rewards with the $t$-wise average and standard deviation across episodes within the batch, i.e., $\{\tilde{r_{t}}\}_{i}=\frac{\{r_{t}\}_{i}-\text{mean}}{\text{std}}$. Then the $t$-th advantage in $i$-th episode can be calculated as:
@@
where $\epsilon$ is a clipping-related hyperparameter for stabilizing training. Finally, the implementation process of the T-ELLM scheme is illustrated in Algorithm 1.
@@
  Collect $T_{m}$ trajectories $[((\mathbf{s}^{m}_{t},\varXi(\mathbb{S}_{t-1}),\mathbb{S}_{t}),(\mathbf{s}^{m}_{t+1},\varXi(\mathbb{S}_{t})))]$ as the training dataset.
1:Add two linear layers (i.e., linear1 and linear2) to GPT-2 model and initialize all parameters $\phi$.
@@
7:              $(\hat{\mathbb{S}}_{t+1},\hat{\varXi}(\mathbb{S}_{t}))=\texttt{linear2}(\text{out}_{t})$.
8:              $\text{loss}_{t}=\texttt{MSE}((\hat{\mathbb{S}}_{t+1},\hat{\varXi}(\mathbb{S}_{t})),({\mathbb{S}}_{t+1},{\varXi}(\mathbb{S}_{t})))$.
@@ V Theoretical Accuracy of the Policy Learned from the Virtual Environment
Our approach separately characterizes the system and statistics parts instead of using a real wireless FL environment. The system part uses resource tools to simulate communication and computing costs, where fixed resource (e.g., bandwidth and transmission power) settings often lead to deterministic state transitions $P(\mathbf{s}^{c}_{t+1}|\mathbf{s}^{c}_{t},(\mathbb{S}_{t},\mathbb{R}_{t}))$, due to the explicit mathematical expressions from (6) to (11). On the other hand, the MDP corresponding to the statistical component is denoted as $\mathcal{M}_{d}=(\mathcal{S}^{m},\mathcal{A}^{m},P^{m},\mathcal{R}^{m},\gamma)$, where, as described in Section IV, the state is defined as $\mathbf{s}^{m}_{t}=\{\mathbf{s}^{m}_{t}(s),\mathbf{s}^{m}_{t}(l)\}\in\mathcal{S}^{m}$. The decision action corresponds to selecting a subset $\mathbb{S}_{t}$ of devices. The transition probability is given by $P^{m}=P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})$, and the reward is defined as the accuracy of the global model, denoted by $\varXi(\mathbb{S}_{t})$. To approximate this statistical MDP, we employ a trained model ^$\hat{\mathcal{M}}_{d}$, whose transition dynamics are captured by the learned function $\hat{P}^{m}$.
@@
Let $P^{m}$ and $\hat{P}^{m}$ denote the transition functions of $\mathcal{M}_{d}$ and ^$\hat{\mathcal{M}}_{d}$, respectively. Suppose the transition model discrepancy is bounded in Total Variation (TV) as:
@@
=$\overset{(a)}{=}$ $\left\|\hat{P}\left(\cdot\mid\mathbf{s}_{t}^{m},\mathbb{S}_{t}\right)-P\left(\cdot\mid\mathbf{s}_{t}^{m},\mathbb{S}_{t}\right)\right\|_{1}$
=$\overset{(b)}{=}$ $2D_{\mathrm{TV}}\left(\hat{P}(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t}),P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})\right)\overset{(c)}{\leq}2\varepsilon.$
@@
where $P_{t}^{\pi}$ and $\hat{P}_{t}^{\pi}$ denote the distributions over $t$-step state-action trajectories induced by policy $\pi$ under transition models $P$ and $\hat{P}$, respectively. The total variation between $K$-step distributions grows at most linearly with the horizon length, as the distributional shift at step $t$ reflects the accumulation of per-step transition errors over the entire trajectory up to time $t$ [52, 53]. In other words, by Corollary 2,
@@
Remark: This result shows that the cumulative advantage error is controllable, growing polynomially with horizon $K$ and linearly with the reward bound and transition error $\varepsilon$. With normalized rewards ($R_{\max}\leq 1$) and a Transformer-based transition model, $\varepsilon$ can typically be reduced to a sufficiently accurate level. Together with proper control of $K$, the deviation between model-based and true advantages remains tightly bound.
@@ VI-A Experimental Setup
TABLE II: Default values of simulation parameters.
@@
Total number of devices $N$ $20$
@@
Number of local iteration $I$ $5$
@@
The total bandwidth of the system $B$ $2$ MHz
Channel coefficient of device $n$ during the $t$-h round $G_{t,n}$ $[10^{-7},10^{-6}]$ dB
The noise power spectral density $N_{0}$ $-174$ dbm/MHz
@@
We consider an image classification task in a wireless FL scenario, as discussed in Section II, where the total number of devices is $N=20$. In the experiments, we use the well-known MNIST dataset for FL training of a Convolutional Neural Network (CNN) model with the cross-entropy loss function. To simulate data heterogeneity across devices, we sample label ratios and dataset sizes from a Dirichlet distribution parameterized by $\alpha$, which controls the degree of non-IIDness. Notably, a smaller $\alpha$ leads to more non-IID data, while a larger $\alpha$ results in more homogeneous data. We set $\alpha=0.2$ as the default to simulate the non-IID data distribution. Other default configurations of simulation parameters are specified in the Table II.
@@
  • $\epsilon$-Greedy Tool[13]: The algorithm is developed based on Oort. Specifically, energy considerations are added, and tools are leveraged to obtain time and energy reference. Moreover, we assume all the device time and energy consumption can be known in advance, so that the $\epsilon$-greedy algorithm can be executed [13].

### v2

@@ III Decoupling of the Optimization Problem
Given the feasible device selection action $\mathbb{S}_{t}$, the performance of global model $\varXi$ is independent of the valid resource allocation action $\mathbb{R}_{t}$, where $\mathbb{R}_{t}$ enable the full participation of devices within the subset $\mathbb{S}_{t}$, i.e.,
@@
$\mathrm{s.t.}\$$\mathbb{S}_{t}\subset\mathcal{N},$ (15)
@@
Remark: Theorem 1 provides the formal basis for decoupling the joint optimization problem. This allows us to deal with the two sub-problems separately and reduces computational complexity by searching over a smaller-dimensional solution space. Accordingly, we can introduce LLM to solve $\mathfrak{P}3$ conveniently, while resorting to a tool-aided solution of $\mathfrak{P}2$. Notably, both sub-problems still involve a calibrated joint design for decision making: For the subproblem $\mathfrak{P}2$, the function $g(\cdot)$ satisfying Assumption 1 can be practically guaranteed by optimization tools. For $\mathfrak{P}2$, the policy itself is required to make optimal device selection decisions under the condition of the resource allocation mechanism $g(\cdot)$.
@@ IV-A Tool-Aided LLM for Efficient Policy Generation
where $\mathbf{s}^{\text{text}}_{t}$ represents the added natural language description. This linguistic representation $\mathbf{o}_{t}$, which will be elaborated later, serves as the input to the LLM. Consequently, it efficiently adapts to diverse state spaces across various scenarios without requiring additional tuning of NN architectures. Let $\mathcal{V}$ represent the LLM’s entire token space. For a given observation $\mathbf{o}_{t}$, the LLM first computes a vector of logits over all tokens in $\mathcal{V}$, i.e., $L_{v}(\mathbf{o_{t}})$. To ensure that the generated action corresponds to a valid device, we define a subset of this token space $\mathcal{I}\subset\mathcal{V}$, as the set of valid action tokens, where each token represents a unique device index $n\in\mathcal{N}$. Therefore, the logits over all valid tokens can be expressed as:
[…]
A softmax function is then applied to this masked logit vector to produce the final probability distribution over the vocabulary, that is, $\pi_{\theta}(\cdot|\mathbf{o}_{t}):=\pi_{\theta}(v|\mathbf{o}_{t})=\frac{\exp(L^{\prime}_{v}(\mathbf{o}_{t}))}{\sum_{u\in\mathbb{V}}\exp(L^{\prime}_{u}(\mathbf{o}_{t}))}$. Then, the LLM generates a text response $\tilde{\mathbf{a}}_{t}$ based on the current language description according to its strategy $\pi_{\theta}$, i.e.,
@@
which is interpreted as the decision action for device selection $\mathbb{S}_{t}$. Upon selecting the subset $\mathbb{S}_{t}$ of devices, the LLM invokes an external resource tool to derive the resource allocation action $\mathbb{R}_{t}$ for the selected devices. Finally, the joint decisions $\mathbf{a}_{t}=[\mathbb{S}_{t},\mathbb{R}_{t}]$ are applied to the wireless FL environment. Thus, the next observation can be obtained by the environment, i.e.,
@@ IV-B Model-Based Virtual Environment
Due to the significant communication overhead associated with online interactions in wireless FL, this section proposes a virtual environment based on an offline model-based approach and tool-assisted computations. Benefiting from the following lemma, the learning of the virtual environment can avoid the extensive use of huge end-to-end expert data.
@@
The state transition of the wireless FL environment can be decomposed into two parts:
@@
By definition, $\mathbf{s}^{c}$ and $\mathbf{s}^{m}$ are conditional independent. Therefore, we can have
@@
$P(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})=P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},(\mathbb{S}_{t},\mathbb{R}_{t}))P(\mathbf{s}^{c}_{t+1}|\mathbf{s}^{c}_{t},(\mathbb{S}_{t},\mathbb{R}_{t})).$ (25)
Recalling that by definition, $\mathbf{s}^{m}$ is $\{|D_{n}|,F_{n}(\omega_{n}),\varXi(\mathbb{S}_{t-1})\}$-related. Then, by Lemma 1, we have the conclusion. ∎
@@
However, for characterizing the statistics part $P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})$, it remains challenging when dynamically selected devices participate in training and aggregation. To address this, we employ a model to approximate the convergence behavior and performance of the global model during FL training. In other words, simulating the statistics part $P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})$ means to predict $\mathbf{s}^{m}_{t+1}$ including $\varXi(\mathbb{S}_{t})$ based on the current observation $\mathbf{s}^{m}_{t}$, device selection action $\mathbb{S}_{t}$ and the historical accuracy $\varXi(\mathbb{S}_{t-1})$. Generally, $\mathbf{s}^{m}_{t}$ can be classified as the state information pertinent to the selected devices $\mathbf{s}^{m}_{t}(s)$ and that of all devices $\mathbf{s}^{m}_{t}(l)$. In particular, $\mathbf{s}^{m}_{t}(s)$ includes post-training local loss $F_{n}(\bm{\omega}_{n})$, the inner-product $\langle\nabla F_{n}(\bm{\omega}_{n}^{t}),\nabla F_{G}(\bm{\omega}_{G}^{t})\rangle$, and the percentage of the same sign $e(\bm{\omega}_{n}^{t},\bm{\omega}_{G}^{t})$ shared between local and global models (i.e., $\bm{\omega}_{n}^{t}$ and $\bm{\omega}_{G}^{t})$) for all selected devices $n\in\mathbb{S}_{t}$. Meanwhile, $\mathbf{s}^{m}_{t}(l)$ consists of the local dataset size $|D_{n}|$ (which remains constant during training) and locally computed loss $l_{n}$ of other non-selected devices. We denote the tuple $((\mathbf{s}^{m}_{t},\varXi(\mathbb{S}_{t-1}),\mathbb{S}_{t}),(\mathbf{s}^{m}_{t+1},\varXi(\mathbb{S}_{t})))$ as one sample of the environment model, while several continuous samples constitute a trajectory. The NN architecture of the simulated model is shown in the right part of Fig. 4, while the details of the training process are shown in Algorithm 1. During training, we optimize the model with a Mean Squared Error (MSE) loss.
Generally, the reward $r_{t}$ can be computed as in (7), except when the resource tool finds no solution exists for $\mathfrak{P}$2, the action $\mathbb{S}_{t}$ will be given a penalty, i.e., $r_{t}=0$. Then, alongside the reward $r_{t}$, the virtual environment can respond to $(\mathbf{s}^{m}_{t},\varXi(\mathbb{S}_{t-1}),\mathbb{S}_{t})$, and accurately yield the next state $(\mathbf{s}^{c}_{t+1},\mathbf{s}^{m}_{t+1})$.
@@ IV-C GRPO-Based LLM Training
The aforementioned virtual environment lays the very foundation for updating the LLM agent in an RL manner. Specifically, we employ GRPO [38] to enhance the inference capability of LLM. As a variant of Proximal Policy Optimization (PPO) [41], GRPO also adopts twin policy models, where an old policy $\pi_{\theta_{\text{old}}}$ is used for sampling actions $\tilde{\mathbf{a}}_{t}$ and accumulating transitional records $(\mathbf{o}_{t},\tilde{\mathbf{a}}_{t},r_{t},\mathbf{o}_{t+1})$ in a replay buffer. Here, $r_{t}$ is the reward in the $t$-th round calculated as in Section IV-B. All samples are described in language, a specific example as shown in Fig. 3. For a batch of $K$-length episodes’ records, where each episode $i$ is denoted as $\{r_{1},\cdots,r_{t},\cdots,r_{K}\}_{i}$, we normalize these rewards with the $t$-wise average and standard deviation across episodes within the batch, i.e., $\{\tilde{r_{t}}\}_{i}=\frac{\{r_{t}\}_{i}-\text{mean}}{\text{std}}$. Then the $t$-th advantage in $i$-th episode can be calculated as:
@@
where $\epsilon$ is a clipping-related hyperparameter for stabilizing training.
Finally, the implementation process of the T-ELLM scheme is illustrated in Algorithm 1.
@@
  Collect $T_{m}$ trajectories $((\mathbf{s}^{m}_{t},\varXi(\mathbb{S}_{t-1}),\mathbb{S}_{t}),(\mathbf{s}^{m}_{t+1},\varXi(\mathbb{S}_{t})))]$ as the training dataset.
1:Add two linear layers (i.e., linear1 and linear2) for embedding to GPT-2 model and initialize all parameters $\phi$.
@@
7:     $({blue}\hat{\mathbf{s}}^{m}_{t+1},\hat{\varXi}(\mathbb{S}_{t}))=\texttt{linear2}(\text{out}_{t})$.
8:     $\text{loss}_{t}=\texttt{MSE}((\hat{\mathbf{s}}^{m}_{t+1},\hat{\varXi}(\mathbb{S}_{t})),({\mathbf{s}}^{m}_{t+1},{\varXi}(\mathbb{S}_{t})))$.
@@ V Theoretical Accuracy of the Policy Learned from the Virtual Environment
Our approach separately characterizes the system and statistics parts instead of using a real wireless FL environment. The system part uses resource tools to simulate communication and computing costs, where fixed resource (e.g., bandwidth and transmission power) settings often lead to deterministic state transitions $P(\mathbf{s}^{c}_{t+1}|\mathbf{s}^{c}_{t},(\mathbb{S}_{t},\mathbb{R}_{t}))$, due to the explicit mathematical expressions in Section II-A2. On the other hand, the MDP corresponding to the statistical component is denoted as $\mathcal{M}_{d}=(\mathcal{S}^{m},\mathcal{A}^{m},P^{m},\mathcal{R}^{m},\gamma)$, where, as described in Section IV, the state is defined as $\mathbf{s}^{m}_{t}=\{\mathbf{s}^{m}_{t}(s),\mathbf{s}^{m}_{t}(l)\}\in\mathcal{S}^{m}$. The decision action corresponds to selecting a subset $\mathbb{S}_{t}$ of devices. The transition probability is given by $P^{m}=P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})$, and the reward is defined as the accuracy of the global model, denoted by $\varXi(\mathbb{S}_{t})$. To approximate this statistical MDP, we employ a trained model $\hat{\mathcal{M}}_{d}$, whose transition dynamics are captured by the learned function $\hat{P}^{m}$.
@@
Let $P^{m}$ and $\hat{P}^{m}$ denote the transition functions of $\mathcal{M}_{d}$ and $\hat{\mathcal{M}}_{d}$, respectively. Suppose the transition model discrepancy is bounded in Total Variation (TV) as:
@@
$\overset{(a)}{=}$ $\left\|\hat{P}\left(\cdot\mid\mathbf{s}_{t}^{m},\mathbb{S}_{t}\right)-P\left(\cdot\mid\mathbf{s}_{t}^{m},\mathbb{S}_{t}\right)\right\|_{1}$
$\overset{(b)}{=}$ $2D_{\mathrm{TV}}\left(\hat{P}(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t}),P(\mathbf{s}^{m}_{t+1}|\mathbf{s}^{m}_{t},\mathbb{S}_{t})\right)\overset{(c)}{\leq}2\varepsilon.$
@@
where $P_{t}^{\pi}$ and $\hat{P}_{t}^{\pi}$ denote the distributions over $t$-step state-action trajectories induced by policy $\pi$ under transition models $P$ and $\hat{P}$, respectively. The total variation between $K$-step distributions grows at most linearly with the horizon length, as the distributional shift at step $t$ reflects the accumulation of per-step transition errors over the entire trajectory up to time $t$ [44]. In other words, by Lemma 3,
@@
Remark: Theorem 2 result shows that the cumulative advantage error is controllable, growing polynomially with horizon $K$ and linearly with the reward bound and transition error $\varepsilon$. Therefore, there are several practical implications. First, the framework is most effective for scenarios with a finite horizon, as the cumulative error may become significant in very long-running tasks. Second, the performance bound is directly tied to the fidelity of the virtual environment. Finally, the reliance on $\tilde{R}_{\max}$ underscores the importance of a well-designed and stable reward function, as volatility or poor scaling can loosen the bound. Fortunately, with normalized rewards ($R_{\max}\leq 1$) and a Transformer-based transition model, $\varepsilon$ can typically be reduced to a sufficiently accurate level. Together with proper control of $K$, the deviation between model-based and true advantages remains tightly bound.
@@ VI-A Experimental Setup
TABLE III: Default values of simulation parameters.
@@
Total number of devices NN 2020
@@
Number of local iteration II 55
@@
The total bandwidth of the system $B$ 22 MHz
Channel coefficient of device $n$ during the $t$-h round $G_{t,n}$ $10^{-7},10^{-6}]$ dB
The noise power spectral density $N_{0}$ 174-174 dbm/MHz
@@
We consider an image classification task in a wireless FL scenario, as discussed in Section II, where the total number of devices is $N=20$. In the experiments, we use the well-known MNIST dataset for FL training of a Convolutional Neural Network (CNN) model with the cross-entropy loss function. To simulate data heterogeneity across devices, we sample label ratios and dataset sizes from a Dirichlet distribution parameterized by $\alpha$, which controls the degree of non-IIDness. Notably, a smaller $\alpha$ leads to more non-IID data, while a larger $\alpha$ results in more homogeneous data. We set $\alpha=0.2$ as the default to simulate the non-IID data distribution. Other default configurations of simulation parameters are specified in the Table III.
@@
  • $\epsilon$-Greedy Tool[7]: A variant of the Oort algorithm. Specifically, energy considerations are added, and tools are leveraged to obtain time and energy reference. Moreover, we assume all the device time and energy consumption can be known in advance, so that the $\epsilon$-greedy algorithm can be executed [7].
@@
  • SAC (Soft Actor-Critic)[10, 46]: SAC is an off-policy actor-critic deep RL algorithm based on the maximum entropy framework[46]. [10] applies SAC to wireless FL scenarios for resource allocation. We implement SAC on top of a transformer decoder-only neural network.